6. Use the method of variation of parameters to find the general solution to the differential equation y - 2y + y = x-le®
Answer #1

Given defferential equation is y 2ylty=xtex. The deedereed equation is y- 2y + y=o. The auxiliary equation m²_amt 1=0 =) (nNow, W(X) = y (8) ฯ) (x) Y, (X) ₂(x) keze 11 1 ex extreer e28 txean xe 2x e 28 +0. . Wic8) = -S x.ed, er at dre ezze - de =-+ or y = Gent cared- xex y = (1 + (2 X) ex & logg- 1) xer, sex logge or, y = (erteilex+ (logx-1) xex Thas the required soluti